### v1

# Jiaxin Wu, Wengyu Zhang, Xiao-Yong Wei⋆†, Qing Li

Abstract
In this paper, we present our methods and results for the Video-To-Text (VTT) task at TRECVid 2024 [1], exploring the capabilities of Vision-Language Models (VLMs) like LLaVA and LLaVA-NeXT-Video in generating natural language descriptions for video content. We investigate the impact of fine-tuning VLMs on VTT datasets to enhance description accuracy, contextual relevance, and linguistic consistency. Our analysis reveals that fine-tuning substantially improves the model’s ability to produce more detailed and domain-aligned text, bridging the gap between generic VLM tasks and the specialized needs of VTT. Experimental results demonstrate that our fine-tuned model outperforms baseline VLMs across various evaluation metrics, underscoring the importance of domain-specific tuning for complex VTT tasks.
1 Video-To-Text (VTT)
The Video-to-Text (VTT) task poses the challenge of generating concise, accurate natural language descriptions for video content, which is a complex vision-language task critical in domains like accessibility, content retrieval, and human-computer interaction. The VTT task requires integrating visual information with language processing. With advancements in Large Language Model(LLM) and Vision Language Model (VLM) like LLaMA [2] and LLaVA [3], researchers have demonstrated the ability of these models to understand visual and textual information.
Therefore, we consider leveraging the power of VLM models in the VTT task for better text description generation. Specifically, we utilize the LLaVA [3] and LLaVA-NeXT-Video [4] model for the VTT task. VLM has been pre-trained on large amounts of visual-textual data and fine-tuned with instructions for specific tasks, such as video understanding, video question answering, and video captioning. We further fine-tune the VLM on a large amount of VTT video-text pairs, aiming to enable the model to specialize in the Video-To-Text task.
2 Method
To generate description text from videos, the following three methods are applied.
Figure 1: Vision Language Model Framework.
2.1 Generating Text Description by Vision Language Models
[…]
$p(\mathbf{X}_{\mathrm{a}}\mid\mathbf{X}_{\mathrm{v}},\mathbf{X}_{\mathrm{q}})=\prod_{i=1}^{L}p(x_{i}\mid\mathbf{X}_{\mathrm{v}},\mathbf{X}_{\mathrm{a},<i},\mathbf{X}_{\mathrm{q},<i}).$
The training process of VLMs typically consists of two stages,
  • Vision-Language Alignment: to align the visual embeddings with LLM’s text embeddings, as LLM’s visual ability acquisition, and
  • Visual Instruction Tuning: to give LLM abilities to complete different kinds of visual taskings, such as video captioning, video short question answering, and video multiple-choice question answering.
LLaVA
We use the LLaVA-v1.5-7b model [^1] https://huggingface.co/liuhaotian/llava-v1.5-7b and follow the official instructions [^2] https://github.com/haotian-liu/LLaVA/tree/main?tab=readme-ov-file#quick-start-with-huggingface to perform the captioning on VTT24 videos. We extract the middle frame of each video as the image input of the model. The text query we used is
"Please write a description of this video frame (around 20-30 words), focusing on
Who, What, Where, and When."
LLaVA-NeXT-Video
We use the LLaVA-NeXT-Video-7B-DPO model [^3] https://huggingface.co/lmms-lab/LLaVA-NeXT-Video-7B-DPO and follow the official instructions [^4] https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA-NeXT-Video.md to perform the captioning on VTT24 videos. The entire video is used as the video input of the model. The prompt we used is:
"Please provide a detailed description of the video, focusing on the main subjects,
their actions, the background scenes."
2.2 Fine-tuning Vision Language Models on VTT Task
The capabilities of vision language models (VLMs) have been investigated on many tasks [6, 7, 8]. However, the vanilla VLMs are typically fine-tuned on specific dataset and instructions, They have noticeable domain gaps on VTT tasks, such as the length of generated text, graininess of description, and wording. Therefore, we decide to fine-tune the VLMs specifically on the VTT dataset for better text description generation. We choose LLaVA [9] as our VLM, and follows official instructions [^5] https://github.com/haotian-liu/LLaVA/blob/main/docs/Finetune_Custom_Data.md to fine-tune the model on VTT dataset. We do not fine-tune the LLaVA-NeXT-Video model since no official fine-tuning code is available.
We collect video-text pairs from VTT16-VTT23 datasets [10]. For each video, we extract frames for each of the 5 frames, resulting in $699683$ frame-text pairs for fine-tuning. The fine-tuning text query we used is
"Please write a description of this video frame (around 20-30 words), focusing on
Who, What, Where, and When."
3 Results analysis
Table 1: Performance comparison among Fine-tuned LLaVA (LV-FT), Vanilla LLaVA (LV) and Vanilla LLaVA-NeXT-Video (LV-V). The best performances are in bold. Rob.: Robustness.
Run Model Task BL $\uparrow$ ME $\uparrow$ CI $\uparrow$ CD $\uparrow$ SP $\uparrow$ S1 $\uparrow$ S2 $\uparrow$ S3 $\uparrow$ S4 $\uparrow$ S5 $\uparrow$
1 LV-FT Main 0.128 0.379 0.712 0.427 0.149 0.448 0.446 0.482 0.446 0.452
3 LV 0.101 0.324 0.637 0.323 0.114 0.432 0.438 0.458 0.418 0.431
4 LV-V 0.027 0.286 0.511 0.015 0.156 0.459 0.447 0.478 0.432 0.451
1 LV-FT Rob. 0.131 0.377 0.715 0.443 0.147 0.444 0.445 0.467 0.444 0.444
3 LV 0.105 0.318 0.634 0.321 0.113 0.432 0.438 0.448 0.414 0.433
4 LV-V 0.027 0.293 0.490 0.016 0.158 0.456 0.441 0.474 0.428 0.445
The evaluation result of 3 proposed methods on VTT24 is shown in Table 1. We report the BLEU (BL), METEOR (ME), CIDEr (CI), CIDEr-D (CD), SPICE (SP) and STS1-5 (S1-S5) scores, aligning with [11].
3.1 The Capability of Pre-trained VLM
We explore the inherent capabilities of the pre-trained Visual Language Model (VLM) without fine-tuning. We focus on evaluating the model’s performance across different tasks when presented with video or frames as inputs.
3.1.1 Video as Input
When video data is directly input to the model, we observe notable performance differences across the three model variants. Table 1 shows that the Vanilla LLaVA-NeXT-Video (LV-V) generally lags in metrics such as BLEU (BL), METEOR (ME), and CIDEr (CI) compared to the fine-tuned and vanilla LLaVA (LV-FT and LV). This suggests that the video variant may not fully leverage sequential data without additional training. Notably, LV-V achieves a higher SPICE (SP) score, indicating a better understanding of semantic relationships in certain scenarios. However, its performance inconsistency implies limited generalization when processing raw video as input.
3.1.2 Frames as Input
In contrast, when input is fed with frames, the vanilla LLaVA (LV) model performs more robustly than the LLaVA-NeXT-Video (LV-V) model across the board. The results in Table 1 reveal that LV achieves higher scores across most metrics, including BLEU, METEOR, and CIDEr. This indicates that frame-based inputs allow LV to capture detailed content more effectively than LV-V, which directly processes video inputs. The consistent advantage of frame-by-frame input suggests that LV better leverages individual frame details to understand and align with ground-truth captions. This approach likely aids the model in capturing nuances that might be lost when processing continuous video sequences as a single input.
3.2 Impact of Fine-tuning
The impact of fine-tuning is evident when comparing LV-FT to the other two models across both main and robustness (Rob.) tasks. LV-FT achieves the highest scores in almost all metrics in the main task, indicating that fine-tuning enhances both syntactic (BLEU, METEOR) and semantic (CIDEr, CIDEr-D) understanding. Specifically, LV-FT’s substantial improvement in CIDEr-D and BLEU scores implies that fine-tuning has refined its ability to capture detailed and relevant content, particularly in challenging scenarios.
Among the 300 queries in VTT24, 171 ($57\%$) showed an improvement in CIDEr scores with the fine-tuned model (LV-FT), compared to the vanilla model (LV). On average, these improved queries achieved a mean gain of 0.356, with the largest observed increase reaching 1.789 in CIDEr score.
LV: A person is parasailing over a beach.
LV-FT: A person is parasailing on a sunny day.
LV: A man is holding a fishing net in the water.
LV-FT: A man in a khaki shirt and a fishing net stands in the water on a sunny day.
LV: A group of children standing in front of a building.
LV-FT: A group of Asian children are standing in front of a building and bowing outside on a sunny day.
LV: A woman with long hair and a black shirt is looking down at the floor.
LV-FT: A woman with long dark hair is lying on the floor and looking at something in front of her.
Figure 2: Case study among Vanilla LLaVA (LV) and Fine-tuned LLaVA (LV-FT)
3.3 Discussions
3.3.1 Fine-tuning Brings VLMs Detailed Description
We discuss the effects of fine-tuning on Vision-Language Models (VLMs), specifically illustrated through the differences in descriptions generated by Vanilla LLaVA (LV) and Fine-Tuned LLaVA (LV-FT) models. Fine-tuning appears to enhance the model’s attention to detail, context, and specificity in its generated descriptions.
In Figure 2, we observe that the LV model typically provides a general description, focusing on the primary subject and basic actions or objects. However, the fine-tuned LV-FT model incorporates additional contextual information and descriptive details, as seen across all examples:
  • In the first image, LV incorrectly describes the ground as "beach". LV-FT, however, refines this by adding "on a sunny day", enriching the setting and potentially suggesting the mood or conditions.
  • Similarly, in the second image, LV’s description "a man holding a fishing net in the water" is accurate yet lacks specificity. LV-FT adds that the man is in a "khaki shirt" and that it’s a "sunny day," thus providing more visual clues.
  • The third image illustrates a significant enhancement in recognizing demographic context: LV mentions "a group of children standing in front of a building," while LV-FT specifies "Asian children bowing outside on a sunny day," adding cultural and situational context, which could improve performance in cultural or geographic datasets.
  • Lastly, in the fourth example, LV’s description is correct but lacks details about her pose. LV-FT adds that she is "lying on the floor and looking at something in front of her," capturing a more complete view of her posture and engagement with the environment.
Overall, these examples demonstrate that fine-tuning enables VLMs to capture richer contextual details and subtle variations in image, which could enhance performance in applications requiring a nuanced understanding of visual content. Fine-tuning not only reinforces the model’s ability to describe the primary elements but also to interpret contextual cues, like clothing color, setting, or implied cultural details. These findings suggest that fine-tuning is an essential step for optimizing VLMs for real-world applications that rely on precise and contextually aware descriptions.
3.3.2 Fine-tuning Aligns VLMs’ Responses with VTT Tasks
(a) VTT16-23
dataset text
(b) Vanilla LLaVA (LV) generated text
(c) Fine-tuned LLaVA (LV-FT) generated text
Figure 3: Comparison among Text Embedding t-SNE Distributions.
Fine-tuning aligns Vision-Language Models (VLMs) with the requirements of VTT tasks by reducing the domain gap, not only improving description detail but also enhancing text style coherence. In Figure 3, the t-SNE visualizations highlight that fine-tuning on VTT16-23 frame-text pairs shifts the distribution of LV-FT’s generated text closer to the target dataset (VTT16-23) than that of Vanilla LLaVA (LV).
This shift indicates several improvements:
  • Domain Alignment: Fine-tuning narrows the gap between generic VLM outputs and the specificity required in VTT, resulting in descriptions that better match the visual and contextual details needed for effective task execution.
  • Linguistic Consistency: LV-FT’s text adopts a style closer to that of VTT16-23 dataset, ensuring descriptions are more cohesive and consistent with domain language patterns.
  • Enhanced Versatility: Compared to LV, LV-FT generates more granular and contextually accurate descriptions. The balanced distribution in LV-FT indicates its adaptability across diverse scenes, improving robustness in VTT applications.
These results highlight the importance of fine-tuning in optimizing VLMs for specialized tasks, enhancing both descriptive quality and consistency in VTT settings.
4 Conclusion
This study demonstrates the effectiveness of fine-tuning Vision-Language Models (VLMs) for the Video-To-Text (VTT) task, highlighting significant improvements in descriptive detail, contextual accuracy, and linguistic alignment. Our experiments show that while pre-trained VLMs exhibit inherent video understanding capabilities, fine-tuning on a VTT-specific dataset enhances their performance across multiple metrics. The fine-tuned LLaVA model (LV-FT) consistently outperforms the vanilla and video-specific models, achieving higher scores in both syntactic and semantic metrics. These findings suggest that VLMs can be optimized for domain-specific tasks by adapting to dataset characteristics, enabling more accurate and context-aware video descriptions. Future work could extend these findings by exploring additional fine-tuning strategies and evaluating their impact on other complex vision-language tasks.
5 Acknowledgments
This research project is supported by the National Natural Science Foundation of China (Grant No.: 62372314).